The volume of a gas at STP is 1.12 x 10-7 ml . Calculate the number of molecules present in the gas .
 

Grade:9

4 Answers

ARUNSOUMYA BASU
7 years ago
At STP, a volume of 22400 mL is occupid by 6.023*1023 molecules of any gas.
So, 1.12 * 10-7mL is occupied by 6.023*1023 *1.12 * 10-7*22400-1 moleules or 3.0115 * 1012 molcules.
Vivek
5 years ago
At S.T.P 1mole=6.023×10^23 molecules which occupy a volume of 22400ml.Let the number of molecules present in 1.12×10^-7ml be `X`.So, 6.023×10^23 ×1.12*10^-7= X × 22400 X=6.023×10^23 ×1.12*10^-7 ÷ 22400 Number of molecules (X)=3.01 ×10^12 molecules
Raghvendra kumar
5 years ago
Here cc mean ml , so here at stp 22400 ml volume is occupied by 6.023 * 10²³ molecules
  1. To calculate no of molecules in 1.12*10¯7  volume of gas
  2. 1.12 * 10¯7​ * 6.023 * 10²³ divided by 22400
  3. No of molecumol is equal to 3.013* 10¹²
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

At S.T.P 1mole=6.023×10^23 molecules which occupy a volume of 22400ml.
Let the number of molecules present in 1.12×10^-7ml be `X`.
So, 6.023×10^23 ×1.12*10^-7= X × 22400
X=6.023×10^23 ×1.12*10^-7 ÷ 22400
Number of molecules (X)=3.01 ×10^12 molecules

Thanks and Regards